# Presenting a STEM Ways of Thinking Framework for Engineering Design-based Physics Problems

### v1

@@
Investigating students’ thinking in classroom tasks, particularly in science and engineering, is essential for improving educational practices and advancing student learning. In this context, the notion of Ways of Thinking (WoT) has gained traction in STEM education, offering a framework to explore how students approach and solve interdisciplinary problems. Building on our earlier studies and contributing to ongoing discussions on WoT frameworks, this paper introduces a new WoT framework—Ways of Thinking in Engineering Design-based Physics (WoT4EDP). WoT4EDP integrates five key elements—design, science, mathematics, metacognitive reflection, and computational thinking—within an undergraduate introductory physics laboratory. This framework offers a novel perspective by emphasizing how these interconnected elements work together to foster deeper learning and holistic problem-solving in Engineering Design (ED)-based projects. We describe the development of WoT4EDP, situate it within the broader landscape of undergraduate STEM education, and provide detailed characterizations of its components. Additionally, we compare WoT4EDP with two contemporary frameworks—Dalal et al. (2021) and English (2023)—to glean insights that enhance its application and promote interdisciplinary thinking. This paper is the first of a two-part series. In the upcoming second part, we will demonstrate the application of the WoT4EDP framework, showcasing how it can be used to analyze student thinking in real-world, ED-based physics projects.
@@ 1 Introduction
Among the several innovations in educational practices is the approach of teaching STEM (science, technology, engineering, and mathematics) by integrating multiple disciplines. [8, 9]. As educators and researchers, we have a responsibility towards preparing students in being “well informed about technical matters and well educated in the STEM subjects” [10, preface.vii]. Students need to be provided both opportunities and guidance to develop a variety of thinking skills. In this context, the notion of ‘Ways of Thinking’ (WoT) has gained traction in STEM education. Few researchers [11, 12, 13, 14] have even proposed WoT frameworks in the context of STEM-based problem solving and Engineering Education Research.
@@
In this paper, we build on our earlier studies [15, 16] on the ‘design-science gap’ [17, 18, 19] in engineering design (ED)   [20] projects in an introductory physics laboratory. In our recent work [21] (currently under review), we had presented an analytic framework to explore and operationalize ‘design thinking’ and ‘science thinking’. We suggested transitioning from the deficit framed notion of design- science ‘gap’ to the asset framed notion of design-science ‘connection’. Continuing on our explorations, and drawing from the works of English [14], Dalal et al. [11], and Slavit et al. [12, 13], we now propose a novel STEM Ways of Thinking framework tailored for Engineering Design (ED)-based physics contexts, where students engage in projects that integrate physics, mathematics, and design thinking to solve real-world-inspired challenges. This framework—referred to as Ways of Thinking for Engineering Design-based Physics (WoT4EDP) throughout this paper—offers a structured approach to investigating and facilitating students’ thinking and learning in undergraduate physics education. WoT4EDP addresses the complexities of integrating design and science, providing a detailed lens for analyzing how students apply physics principles, utilize mathematical reasoning, employ computational tools, and engage in reflective thinking as they iterate on their design to develop practical, effective solutions.
In this study, we have two specific research goals (RGs).
@@
In line with the definitions proposed by Dalal et al. [11, p.109] and English [14, p.1220], we define ‘Ways of Thinking’ in the context of ED-based projects in physics as the approaches students adopt to intentionally think, act, and engage throughout their design projects, guiding their decision-making processes. These ways of thinking constitute a comprehensive framework for exploring, analyzing, evaluating, reflecting, and solving complex, interdisciplinary challenges with a focus on practical solutions. In this context, ways of thinking illuminate how students approach engineering design problems, apply physics concepts, make decisions, and engage in iterative processes to develop conceptual solutions and deepen their understanding. Through this study, we aim to introduce a Ways of Thinking framework into Physics Education Research, and in the process extend the discussion on development and significance of such frameworks.
This paper is the first of a two-part series, with this paper focusing on the conceptual development and comparative analysis of the WoT4EDP framework. In the forthcoming article, extending an earlier study [22], we will apply the framework to new data to gain insights into how students approach and solve engineering design (ED) problems.
@@ 2.1 The National Research Council report - 2013
The National Research Council, in its 2013 report Adapting to a Changing World: Challenges and Opportunities in Undergraduate Physics Education highlights the unique role of physics in shaping students’ minds and society by noting:
@@ 2.5 Why a new SWoT Framework?
Figure 1: The proposed Ways of Thinking for Engineering Design-based Physics (WoT4EDP) framework. The double-headed arrows illustrate the interconnectedness of the Ways of Thinking. A non-exhaustive list of pedagogical goals that the framework can support is presented. The concept of ‘Learning Innovation’ is adapted from English [14].
@@
In 2022 the lead author of the current study embarked on an exploratory study [15] into what we termed as the “Types of Thinking” students engage in within an engineering design (ED)-based laboratory task. An ED problem [20] is client-driven and goal-oriented, set in an authentic context to engage students. It includes constraints such as cost, time, materials, familiar resources, and tools. The solution is a tangible product or process. Another important characteristic of ED problems is that they have multiple solution paths, requiring teamwork to creatively balance constraints and resources [25, p.60, 61].
In 2023 we focused on characterizing the design-science gap in students’ thinking in an ED-task [16]. Building on this study, and drawing inspiration from the ongoing studies on SWoT, we transitioned to using the term ‘STEM Ways of Thinking’ [22], in which we investigated and compared student-groups’ thinking in ‘instructor-assigned’ and ‘student-generated’ ED-problems. To understand deeply how students engage in design and science thinking [21], we developed and applied a detailed analytic framework by invoking the Gioia framework [26] from the qualitative methods literature. Having demonstrated a detailed characterization of students’ SWoT in ED-based tasks, we were convinced that there is space for a new framework - WoT4EDP. Given this motivation, we also explored if WoT4EDP can address some of the gaps in the existing frameworks, and more fundamentally, if WoT4EDP can indeed advance the SWoT discourse.
A key motivation for developing a new framework stems from the distinct context of our study—an engineering design task in a laboratory of an introductory calculus-based physics course for engineering. We had noticed in this context that students naturally interweave various thinking processes, such as design and science thinking. Existing frameworks provide valuable perspectives, yet our observations suggest a need for a model that captures the complex interplay of these thinking modes in a setting that uniquely combines engineering design with physics problem-solving. An important novelty in our framework is the inclusion of ‘computational thinking’, which is absent in the other existing frameworks.
@@
We also recognize an overlap with contemporary frameworks, given that ours is an integrated STEM context. In particular, we will discuss how WoT4EDP compares with the frameworks of Dalal et al. and English in extensive detail in the forthcoming sections. We believe that WoT4EDP is a simple and practical framework for integrated STEM physics educators. Building on prior research, we present a refined SWoT framework, highlighting its features and practical applications for physics educators.
@@ 2.6 Guiding Perspectives
Consistent with our perspective that there can be no one-size-fits-all WoT framework - particularly in educational contexts - and that human understanding is enriched by having diverse viewpoints, we found Weiss’ approach of viewing “evaluation as enlightenment” [27, p.223] to be particularly valuable. In light of the complexities inherent in educational settings, we also resonate with Scriven’s assertion that “the most complete picture of reality can be constructed through the use of multiple, different perspectives” [27, p.222].
@@ 3 Presenting the WoT4EDP Framework
Table 1: A list of earlier qualitative studies that contributed to the development of the WoT4EDP framework, all conducted within an introductory physics laboratory course spanning 13 to 14 weeks in a semester. Note: The lead author was a co-author on the Fall 2022 study [28].
@@
Transcripts of group discussions and final written reports of 14 student-groups To characterize student-groups’ ‘design thinking’ and ‘science thinking’. To explore the connections between student-groups’ design and science thinking. To investigate the influence of scaffolds on student-groups’ design and science thinking. [21] (currently under review)
@@ 3.1 Context
Figure 2: Engineering Design problem (reproduced from [21]) provided to the students in Fall 2021. Wikipedia images were provided, but are not reproduced here. See Table 1.
@@ 3.2 Engineering Design based Labs
In parallel with the hands-on and simulation-based lab tasks, student groups of two or three engaged with an ED problem. Seven essential characteristics of ED, as outlined by Capobianco et al. [25], guided the development of our ED problem: client-driven and goal-oriented, authentic context, constraints, use of familiar materials, resources, and tools, solution is an artifact or process, multiplicity of solutions, and teamwork. In the context of our study, we adopt the definition of ED as “a recursive activity that results in artifacts -- physical or virtual -- as well as processes” [20, p.348]. This includes stages such as: problem scoping and information gathering, idea generation, project realization, communication and documentation of performance results, and optimization. Similar terms such as “ill-structured” [35, p.444], “wicked” [36], and “ill-defined” [37, p.226] problems also find mention in the literature, and, for our purposes, we shall not delve into any subtle differences that may exist among them. One caveat is that our students did not develop an actual physical setup due to time and resource constraints - typical aspects which may be obvious to those who teach large enrollment courses. We, therefore, focused on how students engaged in the ED process to progress to the solution.
@@
Figure 3: A slightly edited version (reproduced from [28]) of the Engineering Design problem given to students in Fall 2022. A related YouTube video was shared. The workstation layout-diagram is not reproduced here. See Table 1.
@@ 3.3 The Road to WoT4EDP
In this section, we outline the process by we identified five key Ways of Thinking (WoT) through the qualitative analysis of student artifacts from Engineering Design-based physics projects, as detailed in our earlier studies (see Table 1).
In our initial study [15] we identified four distinct ‘Types of Thinking’: design, science, mathematical, and metacognitive, using qualitative coding of student work. Building on this foundation, we [28] incorporated computational thinking, leveraging student data from a similar ED-based laboratory sequence to further refine the thinking framework. Further, [16], we investigated the ‘design–science gap’ observed in student projects, providing insight into how these modes of thinking interact. Inspired by ongoing discourse on thinking frameworks [11, 24, 13, 14], we adopted the term ‘Ways of Thinking’ in his next study [22], aligning our framework with contemporary educational research. These efforts culminated in an article (under review) on the ‘design–science connection’ [21], where an elaborate coding scheme was developed to characterize and contextualize student thinking in ED-based physics projects.
@@ 3.4 The Five WoT4EDP Elements
DBT Economic aspects, safety aspects, criteria and constraints, identify stakeholders, contraption dimensions, contraption mechanism, design limitations, feasibility study. [20, 21] Design thinking helps students develop problem-solving skills, creativity, empathy, collaboration, iterative thinking, and multiple perspectives. [47] Incorporate real-world, open-ended problems to encourage creative and user-centered solutions. Use iterative prototyping and reflective feedback sessions to help students refine ideas and increase empathy. Collaborate across disciplines to stimulate innovation. A focus on failure as a learning opportunity strengthens resilience. The ‘design thinking competency model’ [48, p.342] may be used as a guide.
@@
Based on our readings [63, p.4], we gather that the term ‘design thinking’ likely originated with Rowe, who used it as the title of his 1987 book [64], in which he aimed to provide a structured account of problem-solving processes in design. Since then, however, the meaning of the term design thinking has undergone several transformations and has drawn varied interpretations.
Haasi and Laakso [63, p.5] describe design thinking as comprising three key elements: (i) practices, including human-centered approaches, thinking by doing, visualizing, integrating multiple approaches, and collaboration; (ii) cognitive approaches, such as abductive reasoning, reflective reasoning, a holistic perspective, and integrative thinking; and (iii) a mindset characterized by exploration, tolerance for ambiguity, optimism, and a future-oriented outlook.
@@
While the debate over the usage of design thinking is thought-provoking, we find merit in the use of the term in our contexts. For operational purposes, one may view design thinking as a problem-solving approach that emphasizes creativity, user-centered solutions, and iterative processes. It involves understanding users’ needs, brainstorming ideas, prototyping, and testing solutions. Widely applied in fields such as engineering and business, design thinking fosters innovation by blending empathy, experimentation, and practical implementation [65, 48]. In the context of school education, Li et al. argue that design thinking should be regarded as a model of thinking to enhance students’ learning [68, p.97]. They cite several studies showing how students can learn through design based activities and also develop design thinking skills [68, p.98]. A valuable resource for educators is the Educational Designer [69], an international open-access e-journal focused on research-based practices in the design, development, and evaluation of educational materials in mathematics, science, engineering, and technology[68, p.101].
@@
The College Board Standards for College Success [49, p.5] outlines five core science practices that shape scientific inquiry. These include: (i) formulating empirically testable questions and predictions, (ii) collecting relevant data to address these questions, (iii) identifying patterns in data through analysis, (iv) constructing scientific explanations using evidence and knowledge, and (v) applying mathematical reasoning to interpret data and solve problems. These practices emphasize critical thinking and methodological rigor in science education.
The Framework for K-12 Science Education [73],  [38, p.xv] identifies three essential dimensions for providing a high-quality science education: Science and Engineering Practice, which engage students in scientific inquiry and problem-solving; Crosscutting Concepts, which include overarching themes such as patterns and systems that apply across disciplines; and Disciplinary Core Ideas, which focus on the key ideas in physical, life, earth, space, and other sciences. Together, these dimensions help students not only understand scientific content but also grasp how it is acquired and the connections that exist between various scientific disciplines.
The Next Generation Science Standards 2013 document also emphasizes the importance of “instructional flexibility” [38, p.xiv], empowering educators to tailor their approaches based on local contexts to enhance student learning. This flexibility is crucial for integrating disciplines such as mathematics and computation, which naturally intersect with scientific inquiry.
The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum [50, p.2,3] report emphasizes promoting physics-based thinking through several key focus areas. (i) Constructing knowledge: Students are encouraged to independently collect, analyze, and interpret real data, building confidence in their ability to generate scientific insights. (ii) Modeling: Developing abstract representations of physical systems is vital, enabling students to predict outcomes, interpret results, and recognize model limitations. (iii) Designing Experiments: Students need to be guided to ask scientific questions, engineer solutions, and troubleshoot, gaining valuable hands-on experience throughout the process. (iv) Developing Technical and Practical Laboratory Skills: Proficiency with laboratory equipment and understanding the appropriate tools for different measurements are key priorities. (v) Analyzing and Visualizing Data: Students are expected to apply statistical methods, quantify uncertainties, and effectively represent findings through data visualization. (vi) Communicating Physics: Students learn to present scientific results in authentic formats, such as reports and presentations, while developing teamwork and ethical communication skills. Together, these focus areas aim to cultivate a comprehensive set of skills essential for scientific inquiry.
@@
The Committee on the Undergraduate Program in Mathematics in its 2015 report, emphasizes the need to foster a “mathematical habits of mind” [52, p.10,11] by developing students’ cognitive abilities and content knowledge. For cognitive goals, the committee recommends: (i) developing students’ thinking and communication skills to reason logically and express mathematical ideas clearly, (ii) linking mathematical theory with interdisciplinary real-world applications, (iii) using technological tools effectively for problem-solving and exploration, and (iv) fostering mathematical independence through open-ended inquiry. As for the content, the committee recommends undergraduate programs should cover core areas such as calculus, linear algebra, data analysis, computing, and mathematical modeling. A similar perspective is articulated in the 2013 report by the Committee on the Mathematical Sciences in 2025 [74], which emphasizes the need to adapt undergraduate education to meet future demands in research, industry, and interdisciplinary applications.

### v3

@@
Investigating students’ thinking in classroom tasks, particularly in science and engineering, is essential for improving educational practices and advancing student learning. In this context, the notion of Ways of Thinking (WoT) has gained traction in STEM education, offering a framework to explore how students approach and solve interdisciplinary problems. Building on our earlier studies and contributing to ongoing discussions on WoT frameworks, this paper introduces a new WoT framework—Ways of Thinking in Engineering Design-based Physics (WoT4EDP). WoT4EDP integrates five key elements—design, science, mathematics, metacognitive reflection, and computational thinking—within an undergraduate introductory physics laboratory. This novel framework highlights how these interconnected elements foster deeper learning and holistic problem-solving in ED-based projects. A key takeaway is that this framework serves as a practical tool for educators and researchers to design, implement, and analyze interdisciplinary STEM activities in physics classrooms. We describe the development of WoT4EDP, situate it within undergraduate STEM education, and characterize its components in detail. Additionally, we compare WoT4EDP with two contemporary frameworks—Dalal et al. (2021) and English (2023)—to glean insights that enhance its application and promote interdisciplinary thinking. This paper is the first of a two-part series. In the upcoming second part, we will demonstrate the application of the WoT4EDP framework, showcasing how it can be used to analyze student thinking in real-world, ED-based physics projects.
@@ 1 Introduction
Among the several innovations in educational practices is the approach of teaching STEM (science, technology, engineering, and mathematics) by integrating multiple disciplines [8, 9]. As educators and researchers, we have a responsibility towards preparing students in being “well informed about technical matters and well educated in the STEM subjects” [10, preface.vii]. Students need to be provided both opportunities and guidance to develop a variety of thinking skills. In this context, the notion of ‘Ways of Thinking’ (WoT) has gained traction in STEM education. A few researchers [11, 12, 13, 14] have even proposed WoT frameworks in the context of STEM-based problem solving and Engineering Education Research. Our goal in this paper is to propose our own WoT framework, designed specifically for Engineering Design (ED)-based physics learning.
Our confidence in presenting a fresh framework stems from a series of prior qualitative studies [15, 16] on the ‘design-science gap’ [17, 18, 19] in engineering design (ED)-based [20] projects in an introductory physics laboratory. Most notably, in our recently accepted manuscript [21], we had presented an elaborate analytic framework to explore, examine, and operationalize ‘design thinking’ and ‘science thinking’, drawing heavily from various qualitative research methodologies. We investigated how students engaged in design thinking, science thinking, mathematical thinking, and metacognitive reflection while solving an instructor-assigned ED problem. On the basis of our analysis, we suggested transitioning from the deficit-framed notion of design-science ‘gap’ to the asset-framed notion of design-science ‘connection’ (see Table 1).
@@
Continuing on our explorations and drawing inspiration from the works of English [14], Dalal et al. [11], and Slavit et al. [12, 13], we now propose a novel STEM Ways of Thinking framework tailored for Engineering Design (ED)-based physics contexts, wherein students engage in projects that integrate physics, mathematics, and design thinking to solve real-world-inspired challenges. This framework—referred to as Ways of Thinking for Engineering Design-based Physics (WoT4EDP) throughout this paper—offers a structured approach to investigating and facilitating students’ thinking and learning in undergraduate physics education. WoT4EDP addresses the complexities of integrating design and science, providing a detailed analytical lens to examine how students apply physics principles, utilize mathematical reasoning, employ computational tools, and engage in reflective thinking as they iterate on their design to develop practical, effective solutions.
In line with the definitions proposed by Dalal et al. [11, p.109] and English [14, p.1220], we define ‘Ways of Thinking’ in the context of ED-based projects in physics as the approaches students adopt to intentionally think, act, and engage throughout their design projects, guiding their decision-making processes. These interconnected ways of thinking constitute a comprehensive framework for exploring, analyzing, evaluating, reflecting, and solving complex, interdisciplinary challenges with a focus on practical solutions. In our context, ‘Ways of Thinking’ illuminate how students approach engineering design problems, apply physics concepts, make decisions, and engage in iterative processes to develop conceptual solutions and deepen their understanding. Through this study, we aim to introduce a Ways of Thinking framework into Physics Education Research (PER), and in the process extend the discussion on development and significance of such frameworks.
While elaborating on the elements of our framework, we draw upon a range of reform documents with particular emphasis on physics education and PER. Based on our experience both as educators and researchers, we are confident our framework, in addition to extending the discourse on WoT frameworks, offers valuable insights for multiple stakeholders: educators, researchers, and students. WoT frameworks depend on educational contexts which, arguably, are inherently complex. It is essential that we remain open to, and learn from, a spectrum of perspectives. To this end, we engage in a detailed comparison with two recent contemporary WoT frameworks. In this study, we have two specific research goals (RGs).
@@
This theoretical paper is the first in a two-part series. Here, we focus on the conceptual development and comparative analysis of the WoT4EDP framework, grounding it in reform documents, relevant research, and our prior studies.
In the follow-up article, extending our earlier study [22], we will apply the framework to new student-generated data to examine how students approach and solve ED problems. By presenting a transparent and rigorous account of our qualitative analysis, we aim to advance discourse on qualitative methodologies in Physics Education Research.
@@ 2.1 The National Research Council report - 2013
The National Research Council (NRC), in its 2013 report Adapting to a Changing World: Challenges and Opportunities in Undergraduate Physics Education highlights the unique role of physics in shaping students’ minds and society by noting:
@@ 2.5 Why a new SWoT Framework?
Figure 1: The proposed Ways of Thinking for Engineering Design-based Physics (WoT4EDP) framework. The double-headed arrows illustrate the interconnectedness of each element with the other four Ways of Thinking. A non-exhaustive list of pedagogical goals that the framework can support is presented. The concept of ‘Learning Innovation’ is adapted from English [14].
@@
In 2022 the lead author of the current study embarked on an exploratory study [15] into what we termed as the “Types of Thinking” that students engage in within an engineering design (ED)-based laboratory task. An ED problem [20] is client-driven and goal-oriented, set in an authentic context to engage students. It includes constraints such as cost, time, materials, familiar resources, and tools. Given that this is a physics course, we structured the problem to facilitate the application of physics concepts. The solution is a tangible product or process. Another important characteristic of ED problems is that they have multiple solution paths, requiring teamwork to creatively balance constraints and resources [25, p.60, 61].
In 2023, we focused on characterizing the design-science gap in students’ thinking within an ED task [16]. Building on this study, in 2024, inspired by ongoing emphasis on SWoT, we transitioned to using the term ‘STEM Ways of Thinking’ [22] and examined how student groups’ thinking compared when solving both instructor-assigned and student-generated ED problems. Our recently accepted manuscript [21] presents an in-depth investigation of the design-science connections in students’ thinking while solving an instructor-assigned ED problem (see Figure 2) in a physics laboratory. To guide our analysis, we developed and applied a rigorous analytic framework, integrating qualitative research approaches such as traditional coding, thematic analysis, and thick description. Given the lack of universally accepted definitions of ‘design thinking’ and ‘science thinking’, a key objective was to characterize these constructs within our context. Having established a detailed characterization of students’ SWoT in ED-based tasks, we saw an opportunity to contribute to the broader discourse on SWoT frameworks, leading to the development of the WoT4EDP framework.
At this juncture, we confronted a critical question: Is the development of a new framework truly necessary, or would modifying existing ones suffice? As we explain in subsequent sections, the frameworks proposed by Dalal et al. and English were originally designed for contexts that differ significantly from ours. The elements of Dalal et al.’s and English’s frameworks, in our view, do not always directly apply to physics classrooms, though valuable connections can be made depending on local contexts. Our aim is to propose a framework closely aligned with ED-based physics courses while maintaining adaptability for broader application in physics education. Furthermore, we sought to ground our framework in reform documents and research articles specific to Physics Education Research (PER), which may not have been a primary consideration of Dalal et al. and English. Notably, neither framework provides detailed, practical guidance on how to apply their models to actual student data, at least at the time of submission of this manuscript. Additionally, neither incorporates computational thinking—a key aspect in our pedagogical interventions. We also emphasize the interconnectedness of the five elements within our framework, a dimension that may not be as thoroughly explored in their works. While we acknowledge the influence of these existing frameworks, our approach extends beyond merely adding to them.
A primary motivator for the development of a new framework is the distinct context of our study: an engineering design task embedded within an introductory, calculus-based physics laboratory for engineering students. In this setting, students naturally engage in multiple modes of thinking, such as design thinking and science thinking, which intersect in complex ways. While existing frameworks offer valuable perspectives, our experience reveals the need for a model that captures the nuanced interplay of these thinking processes within the integrated context of engineering design and physics problem-solving. We aimed to provide a practical tool which physics educators can employ for their teaching and which students can leverage to direct their learning. In this paper, we will clearly define each element of our framework, grounded in reform documents that are relevant to physics education. The application of our framework will be addressed in a follow-up paper.
@@
Finally, we recognize and acknowledge an overlap of our framework with contemporary frameworks, given that ours is an integrated STEM context. In particular, we will discuss how WoT4EDP compares with the frameworks of Dalal et al. and English in extensive detail in the forthcoming sections. Ultimately, we contend that WoT4EDP represents a simple yet practical framework for integrated STEM physics educators. Building on prior research, we present a refined SWoT framework, highlighting its features and practical applications for physics educators.
@@ 2.6 Guiding Perspectives
In addressing the question, “Why Are You Doing This Study?” [26, ch.2], we were guided by Maxwell’s Qualitative Research Design (2013). We strove earnestly to align our study with the eight key markers of quality in qualitative research outlined by Sarah Tracy: (a) worthy topic, (b) rich rigor, (c) sincerity, (d) credibility, (e) resonance, (f) significant contribution, (g) ethics, and (h) meaningful coherence [27, p.840]. To ensure ‘reliability’ and ‘validity’ [28, 27, 29, 30] of our framework development process, we anchored our approach in widely recognized reform documents published by organizations such as the National Research Council (NRC), the American Association of Physics Teachers (AAPT), the Mathematical Association of America (MAA), the College Board, and others, along with contemporary research articles. These sources provided a solid foundation for identifying and defining the five elements of our framework, shaping its conceptualization in a manner relevant to Physics Education and our specific educational context [26, ch.3].
Regarding the definitions of the five elements—design-based, science-based, mathematics-based, metacognitive reflection, and computational thinking—we recognize that there is no universally agreed-upon definition for such broad and complex concepts. In this context, we found reassurance in the philosophical view that “there is no possibility of attaining a single, ’correct’ understanding of the world” [31]. In her evocative essay The Role of the Researcher in Qualitative Psychology Kiegelmann argues that researcher’s perspective has an important influence on any qualitative study. She aptly quotes Putnam’s succinct words that there is no ultimate “God’s eye view” [31, p.15]. (Kiegelmann, 2002, p. 15). Accordingly, we advocate for a flexible approach that allows these concepts to be adapted to specific educational contexts. These elements are conceptual constructs rather than fixed, measurable entities, and their interpretation may vary depending on the learning environment and objectives. This inherent flexibility enhances the framework’s capacity to address the complexities of real-world educational settings.
@@
Consistent with our perspective that there can be no one-size-fits-all WoT framework - particularly in educational contexts - and that human understanding is enriched by having diverse viewpoints, we found Weiss’ approach of viewing “evaluation as enlightenment” [32, p.223] to be particularly valuable. In light of the complexities inherent in educational settings, we also resonate with Scriven’s assertion that “the most complete picture of reality can be constructed through the use of multiple, different perspectives” [32, p.222].
For the comparison of frameworks, we selected the works of English (2023) [14] and Dalal et al. (2021) [11]. To our knowledge, these are the only available frameworks that have some connection with our context. Additionally, both frameworks provided valuable insights that enriched our understanding of the elements of our framework, making their inclusion both relevant and constructive.
@@ 3 Presenting the WoT4EDP Framework
Table 1: A list of earlier qualitative studies that contributed to the development of the WoT4EDP framework, all conducted within an introductory physics laboratory course spanning 13 to 14 weeks in a semester. Note: The lead author was a co-author in the Fall 2022 study [33].
@@
Transcripts of group discussions and final written reports of 14 student-groups To characterize student-groups’ ‘design thinking’ and ‘science thinking’. To explore the connections between student-groups’ design and science thinking. To investigate the influence of scaffolds on student-groups’ design and science thinking [21] (manuscript accepted for publication).
@@ 3.1 Context
Figure 2: ED problem (reproduced from [21]) provided to students in Fall 2021. Wikipedia images were provided, but are not reproduced here. Students tackled this problem alongside scaffolded physics experiments in the laboratory, applying relevant fundamental physics principles, namely: (linear) momentum, energy, and angular momentum [34]. See [21] and Table 1 for details.
@@ 3.2 Engineering Design based Labs
In parallel with the hands-on and simulation-based lab tasks, student groups of two or three engaged with an ED problem. Seven essential characteristics of ED, as outlined by Capobianco et al. [25], guided the development of our ED problem: client-driven and goal-oriented, authentic context, constraints, use of familiar materials, resources, and tools, solution is an artifact or process, multiplicity of solutions, and teamwork. In the context of our study, we adopt the definition of ED as “a recursive activity that results in artifacts -- physical or virtual -- as well as processes” [20, p.348]. This includes stages such as: problem scoping and information gathering, idea generation, project realization, communication and documentation of performance results, and optimization. Similar terms such as “ill-structured” [40, p.444], “wicked” [41], and “ill-defined” [42, p.226] problems also find mention in the literature, and, for our purposes, we shall not delve into any subtle differences that may exist among them. One caveat is that our students did not develop an actual physical setup due to time and resource constraints - typical aspects which may be obvious to those who teach large enrollment courses. We, therefore, focused on how students engaged in the ED process to progress to the solution.
@@
Figure 3: A slightly edited version (reproduced from [33]) of the ED problem given to students in Fall 2022. A related YouTube video was shared. The workstation layout-diagram is not reproduced here. Students tackled this problem alongside scaffolded physics experiments in the laboratory, applying relevant physics concepts such as momentum, force, vectors, etc. See [33] and Table 1 for details.
@@ 3.3 The Road to WoT4EDP
In this section, we outline the process by which we identified five key Ways of Thinking (WoT) through the qualitative analysis of student artifacts from Engineering Design-based physics projects, as detailed in our earlier studies (see Table 1).
In our initial study [15] we identified four distinct ‘Types of Thinking’: design, science, mathematical, and metacognitive, using qualitative coding of student work. Building on this foundation, we [33] incorporated computational thinking, leveraging student data from a similar ED-based laboratory sequence to further refine the thinking framework. Further, [16], we investigated the ‘design–science gap’ observed in student projects, providing insight into how these modes of thinking interact. Inspired by ongoing discourse on thinking frameworks [11, 24, 13, 14], we adopted the term ‘Ways of Thinking’ in our next study [22], aligning our framework with contemporary educational research. These efforts culminated in an article (under review) on the ‘design–science connection’ [21], where an elaborate coding scheme was developed to characterize and contextualize student thinking in ED-based physics projects.
@@ 3.4 The Five WoT4EDP Elements
DBT Economic aspects, safety aspects, criteria and constraints, identify stakeholders, contraption dimensions, contraption mechanism, design limitations, feasibility study. [20, 21] Design thinking helps students develop problem-solving skills, creativity, empathy, collaboration, iterative thinking, and multiple perspectives. [53] Incorporate real-world, open-ended problems to encourage creative and user-centered solutions. Use iterative prototyping and reflective feedback sessions to help students refine ideas and increase empathy. Collaborate across disciplines to stimulate innovation. A focus on failure as a learning opportunity strengthens resilience. The ‘design thinking competency model’ [54, p.342] may be used as a guide.
@@
Based on our readings [69, p.4], we gather that the term ‘design thinking’ likely originated with Rowe, who used it as the title of his 1987 book [70], in which he aimed to provide a structured account of problem-solving processes in design. Since then, however, the meaning of the term design thinking has undergone several transformations and has drawn varied interpretations.
Haasi and Laakso [69, p.5] describe design thinking as comprising three key elements: (i) practices, including human-centered approaches, thinking by doing, visualizing, integrating multiple approaches, and collaboration; (ii) cognitive approaches, such as abductive reasoning, reflective reasoning, a holistic perspective, and integrative thinking; and (iii) a mindset characterized by exploration, tolerance for ambiguity, optimism, and a future-oriented outlook.
@@
While the debate over the usage of design thinking is thought-provoking, we find merit in the use of the term in our contexts. For operational purposes, one may view design thinking as a problem-solving approach that emphasizes creativity, user-centered solutions, and iterative processes. It involves understanding users’ needs, brainstorming ideas, prototyping, and testing solutions. Widely applied in fields such as engineering and business, design thinking fosters innovation by blending empathy, experimentation, and practical implementation [71, 54]. In the context of school education, Li et al. argue that design thinking should be regarded as a model of thinking to enhance students’ learning [74, p.97]. They cite several studies showing how students can learn through design based activities and also develop design thinking skills [74, p.98]. A valuable resource for educators is the Educational Designer [75], an international open-access e-journal focused on research-based practices in the design, development, and evaluation of educational materials in mathematics, science, engineering, and technology[74, p.101].
@@
The College Board Standards for College Success [55, p.5] outlines five core science practices that shape scientific inquiry. These include: (i) formulating empirically testable questions and predictions, (ii) collecting relevant data to address these questions, (iii) identifying patterns in data through analysis, (iv) constructing scientific explanations using evidence and knowledge, and (v) applying mathematical reasoning to interpret data and solve problems. These practices emphasize critical thinking and methodological rigor in science education.
The Framework for K-12 Science Education [79],  [43, p.xv] identifies three essential dimensions for providing a high-quality science education: Science and Engineering Practice, which engage students in scientific inquiry and problem-solving; Crosscutting Concepts, which include overarching themes such as patterns and systems that apply across disciplines; and Disciplinary Core Ideas, which focus on the key ideas in physical, life, earth, space, and other sciences. Together, these dimensions help students not only understand scientific content but also grasp how it is acquired and the connections that exist between various scientific disciplines.
The Next Generation Science Standards 2013 document also emphasizes the importance of “instructional flexibility” [43, p.xiv], empowering educators to tailor their approaches based on local contexts to enhance student learning. This flexibility is crucial for integrating disciplines such as mathematics and computation, which naturally intersect with scientific inquiry.
The AAPT Recommendations for the Undergraduate Physics Laboratory Curriculum [56, p.2,3] report, published by the American Association of Physics Teachers (AAPT), emphasizes promoting physics-based thinking through several key focus areas. (i) Constructing knowledge: Students are encouraged to independently collect, analyze, and interpret real data, building confidence in their ability to generate scientific insights. (ii) Modeling: Developing abstract representations of physical systems is vital, enabling students to predict outcomes, interpret results, and recognize model limitations. (iii) Designing Experiments: Students need to be guided to ask scientific questions, engineer solutions, and troubleshoot, gaining valuable hands-on experience throughout the process. (iv) Developing Technical and Practical Laboratory Skills: Proficiency with laboratory equipment and understanding the appropriate tools for different measurements are key priorities. (v) Analyzing and Visualizing Data: Students are expected to apply statistical methods, quantify uncertainties, and effectively represent findings through data visualization. (vi) Communicating Physics: Students learn to present scientific results in authentic formats, such as reports and presentations, while developing teamwork and ethical communication skills. Together, these focus areas aim to cultivate a comprehensive set of skills essential for scientific inquiry.
@@
The Committee on the Undergraduate Program in Mathematics in its 2015 report, emphasizes the need to foster a “mathematical habits of mind” [58, p.10,11] by developing students’ cognitive abilities and content knowledge. For cognitive goals, the committee recommends: (i) developing students’ thinking and communication skills to reason logically and express mathematical ideas clearly, (ii) linking mathematical theory with interdisciplinary real-world applications, (iii) using technological tools effectively for problem-solving and exploration, and (iv) fostering mathematical independence through open-ended inquiry. As for the content, the committee recommends undergraduate programs should cover core areas such as calculus, linear algebra, data analysis, computing, and mathematical modeling. A similar perspective is articulated in the 2013 report by the Committee on the Mathematical Sciences in 2025 [80], which emphasizes the need to adapt undergraduate education to meet future demands in research, industry, and interdisciplinary applications.